News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 27 July 2001 11.30am EDT

During the last week the observing schedule was replanned to include two observations of the dwarf nova WZ Sge that was accepted as a Target Of Opportunity (TOO) on 24 July following its detection in outburst. The observation of 3C 334 was impacted and will be rescheduled in a future load.

The first insertion and retraction of the LETG using the new OTG software patch was monitored in real-time on July 25 and seen to complete nominally. All grating operations will now be conducted using the new patch. The observations of the dark moon were also monitored on July 25 and were completed as expected.

The replanned schedule with the observations of WZ Sge and the observations for the coming week shown below.

WZ Sge (TOO)            HRC-S/LETG   Jul 26
G21.5-09                HRC-S        Jul 27
Radiation Belts
M32                     ACIS-S
Cygnus Loop             ACIS-S       Jul 28
NGC 6240                ACIS-S       Jul 29
WZ Sge (TOO)            ACIS-S
DQ Herculis             ACIS-S
2MASSJ125807+2329       ACIS-I
2MASSJ130005+1632       ACIS-I
Radiation Belts                      Jul 30
M 16                    ACIS-I
KES 79                  ACIS-I       Jul 31
NGC 4621                ACIS-S
Cir X-1                 ACIS-S/HETG  Aug 1
CI AQL                  ACIS-S
Radiation Belts
PKS 2153-69             ACIS-S       Aug 2
2S0921-630              ACIS-S/HETG
BRI 0401-1711           ACIS-S       Aug 3
2MASSJ022150+1327       ACIS-I
M15                     HRC-I
LBQS 2201-1834          ACIS-S
A262                    ACIS-S
RGH80                   ACIS-I       Aug 4
Radiation Belts
GRS 1915+105            ACIS-S/HETG
NGC 4494                ACIS-S       Aug 5  
Cir X-1                 ACIS-S/HETG
V 1221+4918             ACIS-I


All spacecraft subsystems continued to operate nominally.

Return to Status Reports